# IK Pegasi and the Double Merger Path to Type Ia Supernovae

### v1

@@ Abstract
Recent Gaia astrometry has revealed thousands of main-sequence ++ white-dwarf binaries (MS++WD) at separations of  ${\sim}\,0.1$$10$ au, including a subset hosting unusually massive ($\gtrsim 0.8$ M) WDs. We argue that s-process enrichment in the non-degenerate companion provides a powerful diagnostic for identifying WDs that formed via mergers in hierarchical triple systems. For a massive WD, standard single-star evolution requires a massive ($\gtrsim 4$ M) progenitor, yet such progenitors produce negligible s-process yields. We define IK Peg–type systems as those exhibiting this mass–yield tension: barium-enhanced companions orbiting WDs too massive to have descended from efficient s-process producers. The well-known system IK Peg exemplifies this class. Applying this framework to published spectroscopic data reveals several additional candidates, and we estimate that a few dozen such systems should exist in the current Gaia sample. If these systems trace inner-binary mergers in primordial triples, they represent observable intermediate stages towards eventual Type Ia supernovae via the double-merger pathway, as predicted by recent population-synthesis models.
@@ I Introduction
Recent Gaia astrometry has revealed a few thousand detached main-sequence ++ white-dwarf (MS++WD) binaries with orbital separations of order  ${\sim}\,1$ au and a broad distribution of eccentricities (Shahaf et al., 2023, 2024). Notably, many of these systems exhibit no significant trend toward circularization, a characteristic typically associated with a past common-envelope episode. A subset of these binaries hosts unusually massive WDs in the 11.31-1.3 M range (Yamaguchi et al., 2024a), although such objects remain rare compared to the general field population (Hallakoun et al., 2024).
@@ III.1 Definition and identification
For decades, IK Peg was the only known post-mass transfer system hosting a massive ($\gtrsim 1$ M) WD and an orbital period on the order of tens of days (e.g. Parsons et al., 2023). Recently, Yamaguchi et al. (2024a) reported five additional such systems, found in the Gaia’s NSS catalog. However, binary systems with massive WD companions and chemical abundance or age inconsistencies exist also with orbital separations of hundreds of days. Motivated by the anomalies observed in IK Peg, we propose that s-process pollution serves as a valuable tracer for identifying post-merger binaries. We suggest the following definition of an IK Peg–type system, as an s-process enriched system hosting a massive $\gtrsim$ 0.8 M WD, with an orbital separation of 0.1–10 au.
@@
Figure 3 displays the maximal theoretical Ba yields across different metallicities and progenitor masses (Karakas and Lugaro, 2016; Karakas et al., 2018) as a function of the expected WD mass, assuming a single-star IFMR (Cunningham et al., 2024). Since any realistic mass transfer or dilution process can only reduce the observable surface enrichment, systems violating this condition are unlikely to descend single-progenitor evolution under any plausible assumptions, provided the companion itself is not an AGB star.
@@
A further challenge in identifying IK Peg–type systems lies in the inherent uncertainty of abundance estimates. Spectral fitting is highly sensitive to the choice of spectral regions, the underlying radiative transfer code, the adopted zero-points (solar abundance scale), and the fitting methodology. These dependencies introduce significant systematic uncertainties (e.g., Blanco-Cuaresma, 2019). However, large-scale population studies provide a means to characterize and calibrate these biases. For instance, benchmarking the pysme-based pipeline of Rekhi et al. (2026) against GALAH DR3 estimates (Buder et al., 2021) reveals a systematic offset of $\lesssim 0.1$ dex (see Appendix B, and see also figure 7 in Yamaguchi et al. 2025a). We thus adopt an enrichment threshold of $0.25$ dex, and define IK Peg–type binaries as systems that exceeds this threshold despite residing in a region of parameter space where theoretical AGB models predict negligible enrichment ($\text{[Ba/Fe]}_{\text{max}}^{\textsc{agb}}\simeq 0$). Consequently, systematic abundance surveys can, at a minimum, distinguish relative enrichment levels and anchor the absolute scale using control samples with independently constrained abundances. Such controls include, for example, MS+WD binaries with massive WD companions that exhibit no s-process enrichment, or single MS stars with comparable mass, metallicity, and magnitude.
@@ VI.1 The double-merger pathway to Type Ia supernovae
SNe Ia, the thermonuclear explosions of WDs, are primary nucleosynthetic sources of heavy elements and serve as “standard candles” for cosmic distance measurement. These events underpinned the discovery of the accelerating expansion of the Universe and dark energy. Yet, despite their fundamental importance, the identity of their progenitor systems remains elusive. While a binary origin involving a WD is well-established, the nature of the companion star and the specific physical mechanism triggering the explosion remain subjects of active debate (see e.g., Maoz et al., 2014; Ruiter and Seitenzahl, 2025, for a review).
@@
The outcome of this channel depends sensitively on the primordial properties of triple systems, including their separation, mass-ratio, and eccentricity distributions, as well as the overall multiplicity fraction. The population-synthesis models of Rajamuthukumar et al. (2023) and Shariat et al. (2025b) demonstrate that hierarchical triples can efficiently produce massive WDs and enable subsequent thermonuclear explosions through this double-merger pathway. Their simulations indicate that this channel can contribute meaningfully to the observed SN Ia rate within current theoretical uncertainties. The IK Peg–type systems identified in this work represent candidate intermediate stages of this evolutionary sequence, offering an empirical window into a channel that has thus far been characterized only theoretically.
@@ VI.2 IK Peg and the SNe Ia rate
We emphasize that the following estimates should be interpreted as conservative upper limits and illustrative context, rather than as strict physical constraints.
@@ VII Outlook
More broadly, the use of chemical enrichment as a tracer of formation history extends well beyond the specific case of MS++WD binaries. Recent discoveries of lithium in Gaia-identified neutron stars underscores the potential of surface abundances to retain signatures of exotic evolutionary channels (El-Badry et al., 2024). Analogously, should some IK Peg–type candidates prove to host neutron stars rather than WDs, their Ba enhancement may point to a history of mass transfer preceding an accretion-induced collapse. Examples for such cases are the massive, enriched systems reported by Yamaguchi et al. (2025a, see Section V.1), or some of the more massive candidates identified in Escorza and De Rosa (2023), Shahaf et al. (2023, 2024) and Yamaguchi et al. (2024a). In this scenario, the detailed feasibility of which remains to be assessed, the s-process enrichment predates the collapse and reflects the progenitor’s evolutionary history rather than the collapse event itself. Thus, chemical tagging emerges as a scalable and powerful tool for reconstructing the evolutionary pathways of compact-object binaries across the mass spectrum.
We thank Jiadong Li, Dan Maoz, Tsevi Mazeh, Johanna Müller-Horn, Ben Pennell, Silvia Toonen, Jaime Villaseñor, and the anonymous referee for helpful discussions. NH acknowledges support from the Planning & Budgeting Committee of the Israeli Council for Higher Education. $S$ and HWR acknowledge support from the European Research Council for the ERC Advanced Grant [101054731]. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC; https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This work is partly based on observations collected at the European Southern Observatory under ESO programme 114.27T5.001.
@@
NH and SS jointly conceived the study, conducted the investigation, and wrote the manuscript, contributing equally to the work. OI led the analysis of the open cluster MS++WD sample (Ironi et al., 2025), while PR led the analysis of the Ba-enriched sample (Rekhi et al., 2024, 2026), including the comparison with GALAH presented in Appendix B. SBA and HWR contributed to the development of the underlying Ba enrichment survey presented in Rekhi et al. (2024, 2026). All authors participated in scientific discussions and provided feedback on the manuscript.
@@ Appendix B Ba abundance estimate uncertainty
To quantify the typical uncertainties in Ba abundance estimates, we compared the values reported in GALAH DR3 (Buder et al., 2021) for the Rekhi et al. (2024) sample with those derived using pysme, following the methodology of Rekhi et al. (2026). The comparison includes 80 systems, and is shown in Figure 5. For each target, all stellar parameters were fixed to the GALAH DR3 values, and only the Ba abundance was varied. We adopted the solar iron abundance used by GALAH DR3, $A(\mathrm{Fe})_{\odot}=7.38$ (Buder et al., 2021).
@@ Appendix C IK Peg–type systems
Table 1 lists potential IK Peg–type systems identified in this work. The component masses and orbital properties of these candidates are illustrated in Figures 7 and 8, superimposed on the full sample from Rekhi et al. (2024, 2026) for comparison. We emphasize that due to inherent selection biases in both the candidate list and the parent samples, robust statistical inferences regarding the differences between these populations cannot currently be drawn.

### v2

# IK Pegasi and the Double-merger Path to Type Ia Supernovae

@@ Abstract
Recent Gaia astrometry has revealed thousands of main-sequence ++ white-dwarf binaries (MS++WD) at separations of  ${\sim}\,0.1$$10$ au, including a subset hosting unusually massive ($\gtrsim 0.8$ M) WDs. We argue that s-process enrichment in the non-degenerate companion provides a powerful diagnostic for identifying WDs that formed via mergers in hierarchical triple systems. For a massive WD, standard single-star evolution requires a massive ($\gtrsim 4$ M) progenitor, yet such progenitors produce negligible s-process yields. We define IK Peg–type systems as those exhibiting this mass–yield tension: barium-enhanced companions orbiting WDs too massive to have descended from efficient s-process producers. The well-known system IK Peg exemplifies this class. Applying this framework to published spectroscopic data reveals several additional candidates, and we estimate that a few dozen such systems should exist in the current Gaia sample. If these systems trace inner-binary mergers in primordial triples, they represent observable intermediate stages toward eventual Type Ia supernovae via the double-merger pathway, as predicted by recent population-synthesis models.
@@ I Introduction
Recent Gaia astrometry has revealed a few thousand detached main-sequence ++ white-dwarf (MS++WD) binaries with orbital separations of order  ${\sim}\,1$ au and a broad distribution of eccentricities (Shahaf et al., 2023, 2024). Notably, many of these systems exhibit no significant trend toward circularization, a characteristic typically associated with a past common-envelope episode. A subset of these binaries host unusually massive WDs in the 11.31-1.3 M range (Yamaguchi et al., 2024a), although such objects remain rare compared to the general field population (Hallakoun et al., 2024).
@@ III.1 Definition and Identification
For decades, IK Peg was the only known post-mass transfer system hosting a massive ($\gtrsim 1$ M) WD and an orbital period on the order of tens of days (e.g. Parsons et al., 2023). Recently, Yamaguchi et al. (2024a) reported five additional such systems, found in the Gaia’s non-single star (NSS) catalog (Gaia Collaboration et al., 2023). However, binary systems with massive WD companions and chemical abundance or age inconsistencies exist also with orbital separations of hundreds of days. Motivated by the anomalies observed in IK Peg, we propose that s-process pollution serves as a valuable tracer for identifying post-merger binaries. We suggest the following definition of an IK Peg–type system, as an s-process enriched system hosting a massive $\gtrsim$ 0.8 M WD, with an orbital separation of 0.1–10 au.
@@
Figure 3 displays the maximal theoretical Ba yields across different metallicities and progenitor masses (Karakas and Lugaro, 2016; Karakas et al., 2018) as a function of the expected WD mass, assuming a single-star IFMR (Cunningham et al., 2024). Since any realistic mass transfer or dilution process can only reduce the observable surface enrichment, systems violating this condition are unlikely to descend from single-progenitor evolution under any plausible assumptions, provided the companion itself is not an AGB star.
@@
A further challenge in identifying IK Peg–type systems lies in the inherent uncertainty of abundance estimates. Spectral fitting is highly sensitive to the choice of spectral regions, the underlying radiative transfer code, the adopted zero-points (solar abundance scale), and the fitting methodology. These dependencies introduce significant systematic uncertainties (e.g., Blanco-Cuaresma, 2019). However, large-scale population studies provide a means to characterize and calibrate these biases. For instance, benchmarking the pysme-based pipeline of Rekhi et al. (2026) against GALAH DR3 estimates (Buder et al., 2021) reveals a systematic offset of $\lesssim 0.1$ dex (see Appendix B, and see also figure 7 in Yamaguchi et al. 2025a). We thus adopt an enrichment threshold of $0.25$ dex, and define IK Peg–type binaries as systems that exceed this threshold despite residing in a region of parameter space where theoretical AGB models predict negligible enrichment ($\text{[Ba/Fe]}_{\text{max}}^{\textsc{agb}}\simeq 0$). Consequently, systematic abundance surveys can, at a minimum, distinguish relative enrichment levels and anchor the absolute scale using control samples with independently constrained abundances. Such controls include, for example, MS+WD binaries with massive WD companions that exhibit no s-process enrichment, or single MS stars with comparable mass, metallicity, and magnitude.
@@ VI.1 The Double-merger Pathway to Type Ia Supernovae
SNe Ia, the thermonuclear explosions of WDs, are primary nucleosynthetic sources of heavy elements and serve as “standard candles” for cosmic distance measurements. These events underpinned the discovery of the accelerating expansion of the Universe and dark energy. Yet, despite their fundamental importance, the nature of their progenitor systems remains elusive. While a binary origin involving a WD is well-established, the nature of the companion star and the specific physical mechanism triggering the explosion remain subjects of active debate (see e.g., Maoz et al., 2014; Liu et al., 2023; Ruiter and Seitenzahl, 2025, for reviews).
@@
The outcome of this channel depends sensitively on the primordial properties of triple systems, including their separation, mass-ratio, and eccentricity distributions, as well as the overall multiplicity fraction. The population-synthesis models of Rajamuthukumar et al. (2023) and Shariat et al. (2025b) demonstrate that hierarchical triples can efficiently produce massive WDs and enable subsequent thermonuclear explosions through this double-merger pathway. Their simulations indicate that this channel can contribute meaningfully to the observed SN Ia rate within current theoretical uncertainties. The IK Peg–type systems identified in this work represent candidate intermediate stages of this evolutionary sequence, offering an empirical window onto a channel that has thus far been characterized only theoretically.
@@ VI.2 IK Peg and the SNe Ia Rate
We emphasize that the following estimates should be interpreted as conservative upper limits and illustrative contexts, rather than as strict physical constraints.
@@ VII Outlook
More broadly, the use of chemical enrichment as a tracer of formation history extends well beyond the specific case of MS++WD binaries. Recent discoveries of lithium in Gaia-identified neutron stars underscore the potential of surface abundances to retain signatures of exotic evolutionary channels (El-Badry et al., 2024). Analogously, should some IK Peg–type candidates prove to host neutron stars rather than WDs, their Ba enhancement may point to a history of mass transfer preceding an accretion-induced collapse. Examples for such cases are the massive, enriched systems reported by Yamaguchi et al. (2025a, see Section V.1), or some of the more massive candidates identified in Escorza and De Rosa (2023), Shahaf et al. (2023, 2024) and Yamaguchi et al. (2024a). In this scenario, the detailed feasibility of which remains to be assessed, the s-process enrichment predates the collapse and reflects the progenitor’s evolutionary history rather than the collapse event itself. Thus, chemical tagging emerges as a scalable and powerful tool for reconstructing the evolutionary pathways of compact-object binaries across the mass spectrum.
We thank Jiadong Li, Dan Maoz, Tsevi Mazeh, Johanna Müller-Horn, Ben Pennell, Silvia Toonen, Jaime Villaseñor, and the anonymous referee for helpful discussions. N.H. acknowledges support from the Planning & Budgeting Committee of the Israeli Council for Higher Education. $S.$ and H.W.R. acknowledge support from the European Research Council for the ERC Advanced Grant (101054731). This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC; https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. This work is partly based on observations collected at the European Southern Observatory under ESO programme 114.27T5.001.
@@
N.H. and S.S. jointly conceived the study, conducted the investigation, and wrote the manuscript, contributing equally to the work. O.I. led the analysis of the open-cluster MS++WD sample (Ironi et al., 2025), while P.R. led the analysis of the Ba-enriched sample (Rekhi et al., 2024, 2026), including the comparison with GALAH presented in Appendix B. S.B.A. and H.W.R. contributed to the development of the underlying Ba enrichment survey presented in Rekhi et al. (2024, 2026). All authors participated in scientific discussions and provided feedback on the manuscript.
@@ Appendix B Ba Abundance Estimate Uncertainty
To quantify the typical uncertainties in the Ba abundance estimates, we compared the values reported in GALAH DR3 (Buder et al., 2021) for the Rekhi et al. (2024) sample with those derived using pysme, following the methodology of Rekhi et al. (2026). The comparison includes 80 systems, and is shown in Figure 5. For each target, all stellar parameters were fixed to the GALAH DR3 values, and only the Ba abundance was varied. We adopted the solar iron abundance used by GALAH DR3, $A(\mathrm{Fe})_{\odot}=7.38$ (Buder et al., 2021).
@@ Appendix C IK Peg–type Systems
Table 1 lists potential IK Peg–type systems identified in this work. The component masses and orbital properties of these candidates are illustrated in Figures 7 and 8, superimposed on the full sample from Rekhi et al. (2024, 2026) for comparison. We emphasize that due to the inherent selection biases in both the candidate list and the parent sample, robust statistical inferences regarding the differences between these populations cannot currently be drawn.